Find the equation of the regression line when x is 77,50, 71,52, 81,94, 96,99, 67 while y is, 82,66, 78, 34,47, 85,99, 99,68
1
Expert's answer
2021-09-21T16:08:01-0400

To find the regression line, the method of least squares is applied. The regression line is given by, "y(hat)= k*x+m" where, "k" is the slope coefficient and "m" is the intercept. In order to find the regression equation we first estimate "k" and "m" using the formula below,

"k=(n*summation(xy)-summation(x)*summation(y))\/(n*summation(x^2)-(summation(x))^2)"

"m=mean(y)-k*mean(x)"

The above data can be summarized as follows,


"n=9,\nsummation(x)=687, \nsummation(y)=658,\nsummation(x^2)=55077,\nsummation(y^2)=51980,\nsummation(x*y)=52578,\nmean(y)=summation(y)\/n=76.333333,\nmean(x)=summation(x)\/n=73.111111"


"k=((9*52578)-(687*658))\/((9*55077)-(687)^2)\n =0.8918(4 decimal places),\nm=76.333333-(0.8918*73.1111111)\n =5.0405(4 decimal places)"


Therefore, equation of the regression line is given as,

"y(hat)=0.8918*x+5.0405."